OpenStudy (anonymous):

To classify the critical point of the plane autonomous system corresponding to the second order equation: x''+ μ(x^2-1)x' ̇+x = 0 in terms of μ where μ is a real valued constant. How to rewrite the equation in matrix notation to change it as a set of first order differential equations. Step by step.

OpenStudy (experimentx):

Hell man ... this is quite difficult

OpenStudy (anonymous):

Not really here is how

OpenStudy (experimentx):

x^2-1 is bugging me

OpenStudy (anonymous):

Define new var y = x' then your eqn is equivalent to

OpenStudy (anonymous):

Wait one more variable u = y' = x''

OpenStudy (anonymous):

Now you have a system x'=y y'=u u +mu(x^2-1)y +x =0

OpenStudy (anonymous):

It is obviously a NON-linear system

OpenStudy (experimentx):

yeah that's the difficulty

OpenStudy (anonymous):

Now, by definition critical point is where the coefficients are all equal to zero

OpenStudy (anonymous):

Sorry, no u needed x'=y y'=-mu(x^2-1) - x

OpenStudy (anonymous):

Critical pont is where y=o and -Mu*(x^2-1) -x =0

OpenStudy (anonymous):

Many thanks Mikael. I will post another question on unstable critical points.

OpenStudy (experimentx):

sorry man ... can't help you at the moment. haven't worked with non linear system.